My problem is;
Solve by factoring and using the principle o zero products
49k^2=36
Thank you very much

Answer by algebrahouse.com(1659)

"My problem is;
Solve by factoring and using the principle o zero products
49k^2=36"

49k^2 - 36 = 0 {subtracted 36 from both sides}

49k^2 - 36 is a difference of two squares, in the form a^2 - b^2,
it factors into the form (a + b)(a - b)

If a^2 is 49k^2, then a = 7k {square root of 49k^2}
If b^2 is 36, then b = 6 {square root of 36}

49k^2 - 36 = 0
(7k + 6)(7k - 6) = 0 {factored into two binomials}
7k + 6 = 0 or 7k - 6 = 0 {set each factor equal to 0}
7k = -6 or 7k = 6 {subtracted 6 and added 6, respectively}
k = -6/7 or k = 6/7 {divided both sides by 7}
www.algebrahouse.com